# A Behaviour and Disease Model of Testing and Isolation

### v1

@@ 3.1.3 The effective reproduction number
$\mathcal{R}(t)=S_{N}(t)\mathcal{R}_{0}^{D}+q_{B}S_{B}(t)\mathcal{R}_{0}^{D}-(1-p_{N}(t))S_{N}(t)\Delta\mathcal{R}-q_{B}p_{B}(t)S_{B}(t)\Delta\mathcal{R}$ (22)
@@
It is important to note that $\mathcal{R}(t)$ is initial condition dependant. For example, when the system starts near equilibrium such that $S_{N}(0)\approx N_{0}$ and $S_{B}(0)\approx B_{0}$, we observe that $\mathcal{R}(0)\approx\mathcal{R}_{0}$.
@@
$\mathcal{R}(0)\approx\mathcal{R}_{0}^{D}-(1-p_{N}(0))\Delta\mathcal{R}$
@@
3.2 The infected steady states
Let $X^{*}$ denote the steady state for compartment $X$. When infection is present, the dynamics of behaviour are governed by
$\frac{\mbox{d}N}{\mbox{d}t}$ $=-\omega(B,T)N+\alpha(N)(B-T)\,.$ (23)
Note that there exists a steady state solution $(N^{*},B^{*})$ with $B^{*}=1-N^{*}$ that depends on the steady state $T^{*}$. That is, $N^{*}$ solves the quadratic
$G(N)=(\omega_{1}-\alpha_{1})N^{2}-(\omega_{2}T^{*}+\omega_{1}+\omega_{3}+\alpha_{2}-\alpha_{1}(1-T^{*}))N+\alpha_{2}(1-T^{*})=0.$
We see that $G(0)=\alpha_{2}(1-T^{*})>0$ and $G(1)=-[(\omega_{2}+\alpha_{1}+\alpha_{2})T^{*}+\omega_{3}]<0$. Hence, given $T^{*}$ there exists a unique $N^{*}\in(0,1)$ with $G(N^{*})=0$.
Now, fix a value for the prevalence of symptomatic cases $O^{*}=I_{N}^{*}+I_{B}^{*}+T^{*}$. Then
$E^{*}$ $=E_{N}^{*}+E_{B}^{*}=\frac{\gamma}{(1-p_{A})\sigma}O^{*}$ (24)
$A^{*}$ $=A_{N}^{*}+A_{B}^{*}=\frac{p_{A}}{(1-p_{A})}O^{*}$ (25)
$R^{*}$ $=R_{N}^{*}+R_{B}^{*}=\frac{\gamma}{\nu(1-p_{A})}O^{*}$ (26)
$S^{*}$ $=S_{N}^{*}+S_{B}^{*}=1-(E^{*}+A^{*}+O^{*}+R^{*})\,.$ (27)
Also, knowing $T^{*}$ fixes $N^{*}$, $B^{*}=1-N^{*}$, and
@@ 3.2 The infected steady states
$\alpha^{*}$ $=\alpha_{1}N^{*}+\alpha_{2}$ (30)
where $I^{*}=I_{N}^{*}+I_{B}^{*}$. Now, setting Equations (4) and (9) to zero gives
Adding Equations (31) and (32) gives
$E_{B}^{*}=\frac{q_{B}\lambda^{*}S^{*}+(\omega^{*}-q_{B}(\omega^{*}+\sigma))E^{*}}{(1-q_{B})(\alpha^{*}+\omega^{*}+\sigma)}\,.$ (33)
Setting Equations (10) and (11) to zero respectively gives
$A_{B}^{*}$ $=\frac{p_{A}\sigma E_{B}^{*}+\omega^{*}A^{*}}{\gamma+\alpha^{*}+\omega^{*}}$ (34)
$I_{B}^{*}$ $=\frac{(1-p_{A})(1-p_{T})\sigma E_{B}^{*}+\omega^{*}I^{*}}{\gamma+\alpha^{*}+\omega^{*}}\,.$ (35)
Next, setting Equation (13) to zero gives
$R_{B}^{*}=\frac{\gamma(A_{B}^{*}+I_{B}^{*}+T^{*})+\omega^{*}R^{*}}{\alpha^{*}+\omega^{*}+\nu}\,.$ (36)
Thus, given $O^{*}$ and $T^{*}$, the steady states are given by
$S_{N}^{*}$ $=S^{*}-S_{B}^{*}$ $S_{B}^{*}$ $=B^{*}-(E_{B}^{*}+A_{B}^{*}+I_{B}^{*}+T^{*}+R_{B}^{*})$
[…]
Finally, setting Equations (3) and (12) to zero gives the simultaneous equations
$T^{*}$ $=\frac{(1-p_{A})p_{T}\sigma}{\gamma}E_{B}^{*}$
$(\lambda^{*}+\omega^{*})S_{N}^{*}$ $=\alpha^{*}S_{B}^{*}+\nu R_{N}^{*}$
which fixes $I^{*}$ and $T^{*}$ (and hence $O^{*}$).
3.3 Stability
To analyse the stability of the system, we begin by expressing the model in a more compact form. Recall that the evolution of the non-behaver population ($N$) follows
$\frac{\mbox{d}N}{\mbox{d}t}=-\omega(B,T)N+\alpha(N)\left(B-T\right)$ (37)
where $B=1-N$. The dynamics of the non-behavers are given by
$\frac{\mbox{d}S_{N}}{\mbox{d}t}$ $=-\lambda S_{N}-\omega(B,T)S_{N}+\alpha(N)S_{B}+\nu R_{N}$ (38)
$\frac{\mbox{d}E_{N}}{\mbox{d}t}$ $=\lambda S_{N}-\sigma E_{N}-\omega(B,T)E_{N}+\alpha(N)E_{B}$
$\frac{\mbox{d}A_{N}}{\mbox{d}t}$ $=p_{A}\sigma E_{N}-\gamma A_{N}-\omega(B,T)A_{N}+\alpha(N)A_{B}$
$\frac{\mbox{d}I_{N}}{\mbox{d}t}$ $=\left(1-p_{A}\right)\sigma E_{N}-\gamma I_{N}-\omega(B,T)I_{N}+\alpha(N)I_{B}$
where $R_{N}=N-S_{N}-E_{N}-A_{N}-I_{N}$. Similarly, the equations governing the behavers are
$\frac{\mbox{d}S_{B}}{\mbox{d}t}$ $=-q_{B}\lambda S_{B}+\omega(B,T)S_{N}-\alpha(N)S_{B}+\nu R_{B}$ (39)
$\frac{\mbox{d}E_{B}}{\mbox{d}t}$ $=q_{B}\lambda S_{B}-\sigma E_{B}+\omega(B,T)E_{N}-\alpha(N)E_{B}$
$\frac{\mbox{d}A_{B}}{\mbox{d}t}$ $=p_{A}\sigma E_{B}-\gamma A_{B}+\omega(B,T)A_{N}-\alpha(N)A_{B}$
$\frac{\mbox{d}I_{B}}{\mbox{d}t}$ $=\left(1-p_{A}\right)\left(1-p_{T}\right)\sigma E_{B}-\gamma I_{B}+\omega(B,T)I_{N}-\alpha(N)I_{B}$
$\frac{\mbox{d}T}{\mbox{d}t}$ $=\left(1-p_{A}\right)p_{T}\sigma E_{B}-\gamma T$
where $R_{B}=B-S_{B}-E_{B}-A_{B}-I_{B}-T$.
To obtain a matrix representation, we define the following matrices and vectors for non-behavers:
and for the behavers:
with
Then the model can be rewritten in matrix form as
$\frac{\mbox{d}N}{\mbox{d}t}$ $=-\omega(B,T)N+\alpha(N)\left(1-N-T\right)$ (40)
$\frac{\mbox{d}T}{\mbox{d}t}$ $=\left(1-p_{A}\right)p_{T}\sigma E_{B}-\gamma T$
with
where $\mathbf{u}_{1}$ is the first standard basis vector, having a first element equal to $1$ and all other elements equal to 00. The Jacobian matrix for this system is structured as
where $\mathbf{J}^{0}$ is a $2\times 2$ matrix given by
$\mathbf{J}^{n0}$ and $\mathbf{J}^{b0}$ are $4\times 2$ matrices given by
$\mathbf{J}^{n0}=\left(\begin{array}[]{cc}\dfrac{\partial\omega}{\partial B}S_{N}+\dfrac{\partial\alpha}{\partial N}S_{B}+\nu&-q_{T}\beta S_{N}-\dfrac{\partial\omega}{\partial T}S_{N}\\[8.5359pt] \dfrac{\partial\omega}{\partial B}E_{N}+\dfrac{\partial\alpha}{\partial N}E_{B}&q_{T}\beta S_{N}-\dfrac{\partial\omega}{\partial T}E_{N}\\[8.5359pt] \dfrac{\partial\omega}{\partial B}A_{N}+\dfrac{\partial\alpha}{\partial N}A_{B}&-\dfrac{\partial\omega}{\partial T}A_{N}\\[8.5359pt] \dfrac{\partial\omega}{\partial B}I_{N}+\dfrac{\partial\alpha}{\partial N}I_{B}&-\dfrac{\partial\omega}{\partial T}I_{N}\end{array}\right),\quad\mathbf{J}^{b0}=\left(\begin{array}[]{cc}-\dfrac{\partial\omega}{\partial B}S_{N}-\dfrac{\partial\alpha}{\partial N}S_{B}-\nu&-q_{B}q_{T}\beta S_{B}+\dfrac{\partial\omega}{\partial T}S_{N}-\nu\\[8.5359pt] -\dfrac{\partial\omega}{\partial B}E_{N}-\dfrac{\partial\alpha}{\partial N}E_{B}&q_{B}q_{T}\beta S_{B}+\dfrac{\partial\omega}{\partial T}E_{N}\\[8.5359pt] -\dfrac{\partial\omega}{\partial B}A_{N}-\dfrac{\partial\alpha}{\partial N}A_{B}&\dfrac{\partial\omega}{\partial T}A_{N}\\[8.5359pt] -\dfrac{\partial\omega}{\partial B}I_{N}-\dfrac{\partial\alpha}{\partial N}I_{B}&\dfrac{\partial\omega}{\partial T}I_{N}\end{array}\right)$
and the other four sub-matrices $\mathbf{J}^{nn}$, $\mathbf{J}^{nb}$, $\mathbf{J}^{bn}$ and $\mathbf{J}^{bb}$ have dimension $4\times 4$, given by
We now assess the stability of the infection-free steady state, where $S_{N}=N_{0}$ and $S_{B}=B_{0}$. The sub-matrices of the Jacobian matrix $J$, evaluated at this equilibrium, are given by
$\mathbf{J}^{0}=\left(\begin{array}[]{cc}\omega_{1}N_{0}+\alpha_{1}B_{0}-\omega(B_{0},0)-\alpha(N_{0})&-\omega_{2}N_{0}-\alpha(N_{0})\\ 0&-\gamma\end{array}\right)$
$\mathbf{J}^{n0}=\left(\begin{array}[]{cc}\omega_{1}N_{0}+\alpha_{1}B_{0}+\nu&-q_{T}\beta N_{0}-\omega_{2}N_{0}\\ 0&q_{T}\beta N_{0}\\ 0&0\\ 0&0\end{array}\right),\quad\mathbf{J}^{b0}=\left(\begin{array}[]{cc}-\omega_{1}N_{0}-\alpha_{1}B_{0}-\nu&-q_{B}q_{T}\beta B_{0}+\omega_{2}N_{0}-\nu\\ 0&q_{B}q_{T}\beta B_{0}\\ 0&0\\ 0&0\end{array}\right)$
and
$\mathbf{J}^{bb}=\left(\begin{array}[]{cccc}-\alpha(N_{0})-\nu&-\nu&-q_{A}q_{B}\beta B_{0}-\nu&-q_{B}\beta B_{0}-\nu\\ 0&-\sigma-\alpha(N_{0})&q_{A}q_{B}\beta B_{0}&q_{B}\beta B_{0}\\ 0&p_{A}\sigma&-\gamma-\alpha(N_{0})&0\\ 0&\left(1-p_{A}\right)\left(1-p_{T}\right)\sigma&0&-\gamma-\alpha(N_{0})\end{array}\right).$
where $B_{0}=1-N_{0}$, $\omega(B_{0},0)=\omega_{1}B_{0}+\omega_{3}$ and $\alpha(N_{0})=\alpha_{1}N_{0}+\alpha_{2}$. The infection-free steady state where $S_{N}=N_{0}$ and $S_{B}=B_{0}$ follows standard results for disease dynamics without a behavioural component and so is stable when $\mathcal{R}_{0}<1$.
@@ 4 Numerical investigations
For the epidemic regime ($\nu=0$), we explore numerically the impact of key model parameters on the dynamics without waning immunity (SEIR-like infection structure), through exploration of the final size, peaks in testing behaviour, and phase diagrams (Section 4.1). For the endemic regime ($\nu>0$), we use a model with waning immunity (SEIRS-like infection structure) to explore the impacts on the steady states and their stability (Section 4.2).
@@
We list the baseline parameters for our numerical investigation in Table 1. We chose the transmission rate, asymptomatic rate, latent period, and immune period to represent a COVID-like illness. Specifically, we set $\beta$ such that the basic reproduction number is $\mathcal{R}_{0}=3.28$[23], we fixed $p_{A}$ and $q_{A}$ at $0.18$ and $0.58$ respectively [24], and fixed $\sigma$ and $\nu$ at $2.5$ [25] and $0.05$ [26] respectively to correspond to a scaled infectious period where $\gamma=1$ representing an average infection period of $10$ days [25]. We selected the behavioural parameters making up $\alpha$ and $\omega$ to replicate waves in testing behaviour observed over multiple SARS-CoV-2 variants in real data [21]. We fixed the reduction in susceptibility due to risk aversion at $q_{B}=0.5$, similarly to previous behaviour and disease models [17]. Lastly, we fixed the parameters $p_{T}$ and $q_{T}$ at $0.9$ and $0.25$ respectively at baseline, and varied these throughout the numerical explorations.
We ran all numerical simulations with initial conditions of $S_{B}(0)=B_{0}$, $I_{N}(0)=10^{-6}$, $S_{N}(0)=1-S_{B}(0)-I_{N}(0)$ and all other compartments empty, where $B_{0}$ is the disease-free equilibrium. The exception to this was when exploring the transitory dynamics without waning immunity, when we instead set $S_{B}(0)=10^{-6}$. This choice was made to ensure the effective reproduction number at time zero was approximately comparable between simulations when investigating the phase diagrams.
We developed and ran code using Python 3.10.12. The code associated with this study is available at https://github.com/Matthew-Ryan1995/BaD_testing_and_isolation.
@@ 4.1.1 Symptomatic final size
Infection peaks show qualitatively similar patterns (Supplementary Figure S2).
Figure 2: Impact of test effectiveness (${p_{T}}$), isolation effectiveness ($1-{q_{T}}$), initial condition of behaviour (${B(0)}$) and the behaviour-free reproduction number (${\mathcal{R}_{0}^{D}}$) on observed symptomatic (${T}$), undetected symptomatic (${I}$), and total symptomatic (${O=I+T}$) final sizes. The colour bars represent the final size as a percentage of the population and are consistent across each row. Note, we vary the behaviour-free reproduction number instead of $\mathcal{R}_{0}$ due to the dependence of $\mathcal{R}_{0}$ on the vertical axes. The dashed grey lines show $\mathcal{R}_{0}=1$ and the black cross indicates the baseline parameter values from Table 1.
@@ 4.1.2 Peaks in testing behaviour
Figure 3: The impact of key model parameters on the peak population willing to test (${B}$) for different behaviour-free reproduction numbers (${\mathcal{R}_{0}^{D}}$). (a) The initial proportion willing to test and isolate ($B(0)$) varies through changing spontaneous uptake ($\omega_{3}$); (b) test effectiveness ($p_{T}$) varies; (c) isolation effectiveness ($1-q_{T}$) varies, and; (d) the social reproduction number ($R_{0}^{B}$) varies through changing the social influence of behaviour ($\omega_{1}$). Note we vary the behaviour-free reproduction number instead of $\mathcal{R}_{0}$ because of the dependence of $\mathcal{R}_{0}$ on the vertical axes. The dashed grey line shows $\mathcal{R}_{0}=1$ and the black cross shows the baseline parameter values from Table 1.

### v2

@@ 3.1.3 The effective reproduction number
$\mathcal{R}(t)=S_{N}(t)\mathcal{R}_{0}^{D}-(1-p_{N}(t))S_{N}(t)\Delta+q_{B}S_{B}(t)\mathcal{R}_{0}^{D}-q_{B}p_{B}(t)S_{B}(t)\Delta$ (24)
@@
The effective reproduction number $\mathcal{R}(t)$ is a time-dependent quantity representing the instantaneous number of new infections that each current infective is producing, accounting for depletion of the susceptible population. When the system starts with initial conditions such that for early times $t$, $S_{N}(t)\approx N_{0}$ and $S_{B}(t)\approx B_{0}$, we note from comparison of (24) and (23) that $\mathcal{R}(t)\approx\mathcal{R}_{0}$.
@@
$\mathcal{R}(0)\approx\mathcal{R}_{0}^{D}-(1-p_{N}(0))\Delta$
@@
3.2 Stability analysis
Here we investigate the stability properties of the system. Our focus is the infection-free equilibrium, for which we derive the conditions ensuring stability. The main result is stated in below, with full details provided in the Appendices.
Proposition 3.1.
Let $E_{B0}$ denote the infection-free equilibrium of the BaD model described in Section 2, with behaviour at its disease-free level $(S_{N},S_{B})=(N_{0},B_{0})$ where $B_{0}=1-N_{0}$ and $N_{0}$ solves Equation (14). $E_{B0}$ is stable if and only if $\mathcal{R}_{0}<1$ where $\mathcal{R}_{0}$ is defined in Section 3.1 (Equation (23)).
Proof.
See Appendix A. ∎
This result shows that, in the absence of infection, the BaD model with testing follows the standard epidemic threshold principle: if $\mathcal{R}_{0}<1$, small outbreaks die out and the system remains at the infection-free equilibrium, while $\mathcal{R}_{0}>1$ makes this equilibrium unstable. In the latter case, the system moves towards the endemic steady states which we describe in the following subsection.
3.3 The endemic steady states
We now characterize the endemic steady states of the model. The key result is summarized in the following proposition, with detailed derivations provided in the Appendices.
Proposition 3.2.
For each compartment $X$, let $X^{*}$ denote its steady state and write $X^{*}=X_{N}^{*}+X_{B}^{*}$. Let the endemic prevalence of symptomatic cases be $O^{*}=I^{*}+T^{*}$. For a given endemic infection prevalence $I^{*}$ and testing prevalence $T^{*}$, the remaining steady states are determined by
$S_{N}^{*}$ $=1-(E^{*}+A^{*}+O^{*}+R^{*})-S_{B}^{*}$ $S_{B}^{*}$ $=B^{*}-(E_{B}^{*}+A_{B}^{*}+I_{B}^{*}+T^{*}+R_{B}^{*})$ (25)
$E_{N}^{*}$ $=\frac{\gamma}{(1-p_{A})\sigma}O^{*}-E_{B}^{*}$ $E_{B}^{*}$ $=\frac{q_{B}\lambda^{*}S^{*}+\big(\omega^{*}-q_{B}(\omega^{*}+\sigma)\big)E^{*}}{(1-q_{B})(\alpha^{*}+\omega^{*}+\sigma)}$ (26)
$A_{N}^{*}$ $=\frac{p_{A}}{1-p_{A}}O^{*}-A_{B}^{*}$ $A_{B}^{*}$ $=\frac{p_{A}\sigma E_{B}^{*}+\omega^{*}A^{*}}{\gamma+\alpha^{*}+\omega^{*}}$ (27)
$I_{N}^{*}$ $=I^{*}-I_{B}^{*}$ $I_{B}^{*}$ $=\frac{(1-p_{A})(1-p_{T})\sigma E_{B}^{*}+\omega^{*}I^{*}}{\gamma+\alpha^{*}+\omega^{*}}$ (28)
$R_{N}^{*}$ $=\frac{\gamma}{\nu(1-p_{A})}O^{*}-R_{B}^{*}$ $R_{B}^{*}$ $=\frac{\gamma(A_{B}^{*}+I_{B}^{*}+T^{*})+\omega^{*}R^{*}}{\alpha^{*}+\omega^{*}+\nu}$ (29)
where the endemic behaviour prevalence $B^{*}=1-N^{*}$ with $N^{*}$ being the unique solution in $(0,1)$ of the quadratic equation
$G(N)=(\omega_{1}-\alpha_{1})N^{2}-(\omega_{2}T^{*}+\omega_{1}+\omega_{3}+\alpha_{2}-\alpha_{1}(1-T^{*}))N+\alpha_{2}(1-T^{*})=0$ (30)
and the equilibrium rates are given by
@@ 3.3 The endemic steady states
$\alpha^{*}$ $=\alpha_{1}N^{*}+\alpha_{2}.$ (33)
Finally, $I^{*}$ and $T^{*}$ are determined by the simultaneous steady-state constraints
$T^{*}=\frac{(1-p_{A})p_{T}\sigma}{\gamma}E_{B}^{*},$ (34)
$(\lambda^{*}+\omega^{*})S_{N}^{*}=\alpha^{*}S_{B}^{*}+\nu R_{N}^{*}.$ (35)
Proof.
See Appendix B. ∎
This result shows that the endemic equilibrium is defined through a system of coupled equations reflecting the feedback loop between infection dynamics and testing behaviour. Due to this complexity, the endemic states cannot be solved analytically in a straightforward way. We therefore explore their properties numerically in Section 4.2.
@@ 4 Numerical investigations
For the epidemic regime ($\nu=0$), we explore numerically the impact of key model parameters on the dynamics without waning immunity (SEIR-like infection structure), through exploration of the final size, peaks in testing behaviour, and phase diagrams (Section˜4.1). For the endemic regime ($\nu>0$), we use a model with waning immunity (SEIRS-like infection structure) to explore the impacts on the steady states and their stability (Section˜4.2).
@@
We list the baseline parameters for our numerical investigation in Table˜1. We chose the transmission rate, asymptomatic rate, latent period, and immune period to represent a COVID-like illness. Specifically, we set $\beta$ such that the basic reproduction number is $\mathcal{R}_{0}=3.28$[23], we fixed $p_{A}$ and $q_{A}$ at $0.18$ and $0.58$ respectively [24], and fixed $\sigma$ and $\nu$ at $2.5$ [25] and $0.05$ [26] respectively to correspond to a scaled infectious period where $\gamma=1$ representing an average infection period of $10$ days [25]. We selected the behavioural parameters making up $\alpha$ and $\omega$ to replicate waves in testing behaviour observed over multiple SARS-CoV-2 variants in real data such as those seen in Figure 1 of Eales et al. [27]. The transient dynamics for the baseline parameter values of Table 1 are found in Supplementary Figure 9. We fixed the reduction in susceptibility due to risk aversion at $q_{B}=0.5$, similarly to previous behaviour and disease models [19]. Lastly, we fixed the parameters $p_{T}$ and $q_{T}$ at $0.9$ and $0.25$ respectively at baseline, and varied these throughout the numerical explorations.
We ran all numerical simulations with initial conditions of $S_{B}(0)=B_{0}$, $I_{N}(0)=10^{-6}$, $S_{N}(0)=1-S_{B}(0)-I_{N}(0)$ and all other compartments empty, where $B_{0}$ is the disease-free equilibrium. The exception to this was when exploring the transitory dynamics without waning immunity, when we instead set $S_{B}(0)=10^{-6}$. This choice was made to ensure the effective reproduction number at time zero was approximately comparable between simulations when investigating the phase diagrams. We developed and ran code using Python 3.10.12, which we have made open-source and publicly available via GitHub at https://github.com/Matthew-Ryan1995/BaD_testing_and_isolation.
@@ 4.1.1 Symptomatic final size
Infection peaks show qualitatively similar patterns (Supplementary Figure 10).
Figure 2: Impact of test effectiveness ($p_{T}$}), isolation effectiveness ($1-{q_{T}}$), initial condition of behaviour (B(0){B(0)}) and the behaviour-free reproduction number ($\mathcal{R}_{0}^{D}$}) on observed symptomatic (T{T}), undetected symptomatic (I{I}), and total symptomatic (O=I+T{O=I+T}) final sizes. The colour bars represent the final size as a percentage of the population and are consistent across each row. Note, we vary the behaviour-free reproduction number instead of $\mathcal{R}_{0}$ due to the dependence of $\mathcal{R}_{0}$ on the vertical axes. The dashed grey lines show $\mathcal{R}_{0}=1$ and the black cross indicates the baseline parameter values from Table 1.
@@ 4.1.2 Peaks in testing behaviour
Figure 3: The impact of key model parameters on the peak population proportion willing to test (B{B}) for different behaviour-free reproduction numbers ($\mathcal{R}_{0}^{D}$}). (a) The initial proportion willing to test and isolate ($B(0)$) varies through changing spontaneous uptake ($\omega_{3}$); (b) test effectiveness ($p_{T}$) varies; (c) isolation effectiveness ($1-q_{T}$) varies, and; (d) the social reproduction number ($R_{0}^{B}$) varies through changing the social influence of behaviour ($\omega_{1}$). Note we vary the behaviour-free reproduction number instead of $\mathcal{R}_{0}$ because of the dependence of $\mathcal{R}_{0}$ on the vertical axes. The dashed grey line shows $\mathcal{R}_{0}=1$ and the black cross shows the baseline parameter values from Table 1.